# Computational Hypergraph Discovery, a Gaussian Process framework for connecting the dots

### v1

@@ 2.4. Algorithm Overview: An Informal Summary
    Selecting a Kernel: We choose a kernel function, such as linear, quadratic, or fully nonlinear kernels (refer to Example 4.4). The kernel selection process is analogous to the subsequent pruning steps, involving the determination of a parameter $\gamma$, regression analysis, and evaluation based on signal-to-noise ratios.
@@
      Regression Analysis: Predict the node’s value using the current set of ancestors, excluding the least active one (i.e., the one contributing the least to the signal). We employ Kernel Ridge Regression with the selected kernel function and parameter $\gamma$ (see Sec. 4.3.3 and 4.3.3).
@@
      Evaluate Removal: Compute the regression signal-to-noise ratio (see Sec. 4.3.4 and 6):
@@
2.5. Refinements.
Throughout this paper, the approach presented in this section is refined and generalized, as discussed in the following subsections.
2.5.1. Complexity Reduction with Kernel PCA Variant.
To improve the efficiency of our approach, we introduce a variant of Kernel PCA [23]. This modification significantly reduces the computational complexity of our proposed method, making it primarily dependent on the number of principal nonlinear components in the underlying kernel matrices rather than the number of data points. See Sec. 4.3.2.
We extend the concept of descendants and ancestors to cover more complex functional dependencies between variables, including implicit ones. This generalization is achieved through a Kernel-based adaptation of Row Echelon Form Reduction, initially designed for affine systems, and leveraging the principles of Kernel Mode Decomposition [30]. See Sec. 4.3.3.
2.5.3. Parameter Selection.
The choice of the parameter $\gamma$ is a critical aspect of our proposed approach. We provide a structured approach for selecting $\gamma$ based on the characteristics of the kernel matrix $K_{s}$. Specifically, when $K_{s}$ is derived from a finite-dimensional feature map $\psi$, we employ the regression residual to determine $\gamma$ as follows: $\gamma=\min_{v}\big{\|}v^{T}\psi(X)-Y\big{\|}_{\mathbb{R}^{N}}^{2}$. This process is detailed in Sec. 6.2.1. Alternatively, when $K_{s}$ is a universal kernel, we select $\gamma$ by maximizing the variance of the eigenvalue histogram of $\gamma\big{(}K_{s}(X,X)+\gamma I\big{)}^{-1}$ (see Sec. 6.2.2).
2.5.4. Ancestor pruning.
The most efficient version of our proposed algorithm (Alg. 2) does not use a threshold of $0.5$ on the signal-to-noise ratio to prune ancestors, but it rather employs an inflection point in the noise-to-signal ratio $\frac{\mathcal{V}(n)}{\mathcal{V}(s)+\mathcal{V}(n)}(q)$ as a function of the number $q$ of ancestors (Fig. 12). To put it simply, after ordering the ancestors in decreasing contribution to the signal, the final number $q$ of ancestors is determined as the maximizer of $\frac{\mathcal{V}(n)}{\mathcal{V}(s)+\mathcal{V}(n)}(q+1)-\frac{\mathcal{V}(n)}{\mathcal{V}(s)+\mathcal{V}(n)}(q)$. We illustrate this algorithm in the following section with an application to the recovery of a mechanical network.
2.6. The Fermi-Pasta-Ulam-Tsingou problem
The Fermi-Pasta-Ulam-Tsingou (FPUT) problem, delineated in [31], is a renowned dynamical system for exploring chaotic systems.
It comprises a set of $N$ masses, with mass $j\in\{0,\ldots,N-1\}$ having a position at equilibrium $p_{j}=jh$ with $h=1/N$. Each mass is tethered to its two adjacent masses by a spring, and an additional force defined by a function $\alpha\,:\,\mathbb{R}\to\mathbb{R}$. The displacement of the mass $x_{j}$ adheres to the equation:
$\ddot{x}_{j}=\frac{c^{2}}{h^{2}}(x_{j+1}+x_{j-1}-2x_{j})\left(1+\alpha(x_{j+1}-x_{j-1})\right)$ (2.12)
Here, $c^{2}$ is a velocity depending on $h$ and the strength of the springs. We set $c=1$ and $N=10$. Furthermore, we use fixed boundary conditions by adding two more masses, with $x_{-1}=x_{N}=0$.
The selection of this model to illustrate our framework is driven by several aspects: its chaotic behavior, established and sparse equations, and the presence of linear/nonlinear equations contingent on the choice of $\alpha$. We simulate the system with two different choices of $\alpha$: a nonlinear case where $\alpha(x)=x^{2}$, and a linear case where $\alpha(x)=0$. Snapshots of the system, encapsulating the position, velocity, and acceleration of each mass, are collected at various time points. A thousand of these snapshots serve as data for our graph discovery algorithm. We use the algorithm from the preceding section, refined with the optimal noise parameter from Sec. 6, the Z-score test, and the signal-to-noise ratio inflection point to select ancestors.
Figure 11. (a) Recovered full graph (b) Recovered subgraph for positions and accelerations.
The resultant graph, depicted in Fig.11.(a), showcases the graph discovery framework’s aptitude in accurately identifying dependencies. Notably, speed information is accurately identified as non-essential and omitted from the ancestors of position and acceleration, underscoring the efficacy of the approach. The recovered dependencies are notably precise, especially considering the model’s nonlinear nature. Fig.11.(b), which omits speed information for clarity, further elucidates the accurate recovery of dependencies. These dependencies are predominantly the simplest and most straightforward possible, with an exception for $\ddot{x}_{1}$ and $\ddot{x}_{8}$, which have respective recovered ancestors $\{\ddot{x}_{0},x_{1},x_{2}\}$ and $\{\ddot{x}_{9},x_{8},x_{7}\}$, instead of $\{x_{0},x_{1},x_{2}\}$ and $\{x_{9},x_{8},x_{7}\}$. However, examining the equations for $\ddot{x}_{0}$ and $\ddot{x}_{9}$ reveals that they contain the information of $(x_{0},x_{1})$ and $(x_{8},x_{9})$ respectively, validating the recovered relationships as alternative formulations of the FPUT equations. It is crucial to note that multiple formulations of the same equations, with varying ancestor relationships, can generally exist.
Figure 12. (a) Noise-to-signal ratio $\frac{\mathcal{V}(n)}{\mathcal{V}(s)+\mathcal{V}(n)}(q)$ as a function of the number $q$ of proposed ancestors. We also plot the $Z$-test quantiles describe in Sec. 6.3 (in the absence of signal, the noise-to-signal ratio should fall within the shaded area with probability $0.9$).(b) Noise-to-signal ratio increments $\frac{\mathcal{V}(n)}{\mathcal{V}(s)+\mathcal{V}(n)}(q)-\frac{\mathcal{V}(n)}{\mathcal{V}(s)+\mathcal{V}(n)}(q-1)$ as a function of the number $q$ of ancestors. The x-axis represents the number $q$ of proposed ancestors for $\ddot{x}_{7}$. The removal of crucial ancestors, such as ancestor $3$, results in the Noise-to-Signal Ratio saturating at a value of $1.0$. This particular removal also leads to a significantly greater increase in the Noise-to-Signal Ratio compared to the removal of other ancestors.
Figure 13. (a) Noise-to-Signal Ratio, denoted as $\frac{\mathcal{V}(n)}{\mathcal{V}(s)+\mathcal{V}(n)}(q)$, with respect to the number of proposed ancestors, represented by $q$. Additionally, we include a visualization of the quantiles derived from the $Z$-test, as described in Section 6.3. Notably, when there is no signal present, the noise-to-signal ratio is expected to fall within the shaded area with a probability of 0.9. (b) Increments in the Noise-to-Signal Ratio, defined as $\frac{\mathcal{V}(n)}{\mathcal{V}(s)+\mathcal{V}(n)}(q)-\frac{\mathcal{V}(n)}{\mathcal{V}(s)+\mathcal{V}(n)}(q-1)$, as a function of the number of ancestors, denoted as $q$. The horizontal axis represents the number of proposed ancestors for $\dot{x}_{0}$. Determining an appropriate stopping point based solely on absolute noise-to-signal ratio levels can be challenging. In contrast, the increments in the noise-to-signal ratio clearly exhibit a discernible maximum, offering a practical point for decision-making.
While performing our experiments, a set threshold of 60% on the noise-to-signal ratio could be used to obtain the desired equations. However, this choice could be seen as arbitrary or guided by our knowledge of the true equations. However, the choice to retain or discard an ancestor becomes clear when examining the noise-to-signal ratio’s evolution against the number of ancestors. In the graph discovery phase, every other node is initially deemed a potential ancestor for a specified node of interest. We then proceed to iteratively remove the node with the least signal contribution. The step resulting in the largest surge in the noise-to-signal ratio is inferred as one eliminating a crucial ancestor, thereby pinpointing the final ancestor set. In the context of the FPUT problem, we observed that, firstly, removing a node essential to the equation of interest causes the noise-to-signal ratio to markedly jump from approximately 25% to 99%. Additionally, we can analyze the increase in the noise-to-signal ratio during the removal process. Two representative trajectories illustrating this are presented in Fig. 12 and 13. Notably, in the final iteration where an essential node was removed, the increase in the noise-to-signal ratio is significantly higher compared to previous removals. This observation indicates that the process halted at the appropriate juncture, as the removal of this last node was not sanctioned. As shown in Fig.13, while solely relying on a fixed threshold to decide when to cease the removals might prove challenging, evaluating the increments in noise-to-signal ratios offers a clear guideline for efficiently and reliably pruning ancestors.
3. Hardness and well-posed formulation of Type 3 problems
@@
Having discussed why we should care about this problem, we will describe why the problem is challenging and why it can even be intractable if not formalized and approached properly.
Figure 14. Computational Hypergraph Discovery with three variables
@@
3.2. Nonidentifiability and implicit dependencies.
Secondly, it is important to note that, even with an infinite amount of data, the exact structure of the hypergraph might not be identifiable. To illustrate this point, let’s consider a problem where we have $N$ samples from a computational graph with variables $x$ and $y$. The task is to determine the direction of functional dependency between $x$ and $y$. Does it go from $x$ to $y$ (represented as $x$$y$$f$ ), or from $y$ to $x$ (represented as $y$$x$$f$ )?
If we refer to Figure 15.(a), we can make a decision because $y$ can only be expressed as a function of $x$. In contrast, if we examine Figure 15.(b), the decision is also straightforward because $x$ can solely be written as a function of $y$. However, if the data mirrors the scenario in Figure 15.(c), it becomes challenging to decide as we can write both $y$ as a function of $x$ and $x$ as a function of $y$. Further complicating matters is the possibility of implicit dependencies between variables. As illustrated in Figure 15.(d), there might be instances where neither $y$ can be derived as a function of $x$, nor $x$ can be represented as a function of $y$.
@@
The field of causal inference and probabilistic modeling, as outlined in [24, 15], seeks to address distinct graph learning problems by introducing auxiliary assumptions. Within noise and probabilistic models [43, 21], it is generally assumed that the data is randomized.
@@
Causal inference methods broadly consist of two approaches: constraint and score-based methods. While constraint-based approaches are asymptotically consistent, they only learn the graph up to an equivalence class [42]. Instead, score-based methods resolve ambiguities in the graph’s edges by evaluating the likelihood of the observed data for each graphical model. For instance, they may assign a higher evidence to $y\to x$ over $x\to y$ if the conditional distribution $x|y$ exhibits less complexity than $y|x$. The complexity of searching over all possible graphs, however, grows super-exponentially with the number of variables. Thus, it is often necessary to use approximate, but more tractable, search-based methods [7, 32] or alternative criteria based on sensitivity analysis [9]. For example, the preference could lean towards $y\to x$ rather than $x\to y$ if $y$ demonstrates less sensitivity to errors or perturbations in $x$. In contrast, our proposed Gaussian Process method and avoids the growth in complexity by performing a guided pruning process that assesses the contribution of each node to the signal, as detailed in Section 2.3. We also emphasize that our method is not limited to learning acyclic graph structures as it can identify feedback loops between variables.
Alternatively, methods for learning probabilistic undirected graphical models, also known as Markov networks, identify the graph structure by assuming the data is randomly drawn from some probability distribution [12]. In this case, edges in the graph (or lack thereof) encode conditional dependencies between the nodes. A common approach learns the graph structure by modeling the data as being drawn from a multivariate Gaussian distribution with a sparse inverse covariance matrix, whose zero entries indicate pairwise conditional independencies [13]. Recently, this approach has been extended using models for non-Gaussian distributions, e.g., in [1, 33], as well as kernel-based conditional independence tests [46]. In this work, we learn functional dependencies rather than causality or probabilistic dependence. We emphasize that we also do not assume the data is randomized or impose strong assumptions, such as additive noise models, in the data-generating process.
@@
We complete this subsection by comparing the hypergraph discovery framework to structure learning for Bayesian networks and structural equation models (SEM). Let $x\in\mathbb{R}^{d}$ be a random variable with probability density function $p$ that follows the autoregressive factorization $p(x)=\prod_{i=1}^{d}p_{i}(x_{i}|x_{1},\dots,x_{i-1})$ given a prescribed variable ordering. Structure learning for Bayesian networks aims to find the ancestors of variable $x_{i}$, often referred to as the set of parents $Pa(i)\subseteq\{1,\dots,i-1\}$, in the sense that $p_{i}(x_{i}|x_{1},\dots,x_{i-1})=p_{i}(x_{i}|x_{Pa(i)})$. Thus, the variable dependence of the conditional density $p_{i}$ is identified by finding the parent set so that $x_{i}$ is conditionally independent of all remaining preceding variables given its parents, i.e., $x_{i}\perp x_{1:i-1\setminus Pa(i)}|x_{Pa(i)}$. Finding ancestors that satisfy this condition requires performing conditional independence tests, which are computationally expensive for general distributions [39]. Alternatively, SEMs assume that each variable $x_{i}$ is drawn as a function of its ancestors with additive noise, i.e, $x_{i}=f(x_{Pa(i)})+\epsilon_{i}$ for some function $f$ and noise $\epsilon$ [32]. For Gaussian noise $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, each marginal conditional distribution in a Bayesian network is given by $p_{i}(x_{i}|x_{1:i-1})\propto\exp(-\frac{1}{2\sigma^{2}}\|x_{i}-f(x_{1:i-1})\|^{2})$. Thus, finding the parents for such a model by maximum likelihood estimation corresponds to finding the parents that minimize the expected mean-squared error $\|x_{i}-f(x_{Pa(i)})\|^{2}.$ We will show that our approach minimizes a related objective in Section 4.3.3, without imposing the strong probabilistic assumptions that are required in SEMs and Bayesian Networks. We also observe that while the graph structure identified in Bayesian networks is influenced by the specific sequence in which variables are arranged (a concept exploited in numerical linear algebra [37, 35] where Schur complementation is equivalent to conditioning GPs and a carefully ordering leads to the accuracy of the Vecchia approximation $p_{i}(x_{i}|x_{1},\dots,x_{i-1})\approx p_{i}(x_{i}|x_{i-k},\dots,x_{i-1})$ [45]), the graph recovered by our approach remains unaffected by any predetermined ordering of those variables.
@@
Let $\mathcal{H}$ be a Reproducing Kernel Hilbert Space (RKHS) of functions mapping $\mathbb{R}^{d}$ to $\mathbb{R}$. Let $\mathcal{F}$ be a closed linear subspace of $\mathcal{H}$ and let $\mathcal{M}$ be a subset of $\mathbb{R}^{d}$ such that $x\in\mathcal{M}$ if and only if $f(x)=0$ for all $f\in\mathcal{F}$.
Given the (possibly noisy and nonrandom) observation of $N$ elements, $X_{1},\ldots,X_{N}$, of $\mathcal{M}$ approximate $\mathcal{M}$.

### v2

@@ 2. Algorithm Overview for Type 3 problems: An Informal Summary
    Selecting a Kernel: We choose a kernel function, such as linear, quadratic, or fully nonlinear kernels (refer to Example 1). The kernel selection process is analogous to the subsequent pruning steps, involving the determination of a parameter $\gamma$, regression analysis, and evaluation based on signal-to-noise ratios.
@@
      Regression Analysis: Predict the node’s value using the current set of ancestors, excluding the least active one (i.e., the one contributing the least to the signal). We employ Kernel Ridge Regression with the selected kernel function and parameter $\gamma$ (see Sec. 5.3.3 and 5.3.3).
@@
      Evaluate Removal: Compute the regression signal-to-noise ratio (see Sec. 5.3.4 and 7):
@@
Figure 7. Formal description of Type 2 problems.
3. Type 2 problems: Formal description and GP-based Computational Graph Completion
3.1. Formal description of Type 2 problems
3.2. Reminder on Computational Graph Completion for Type 2 problems
Within the context of Sec. 3.1, the proposed GP solution to Type 2 problems is to simply replace unknown functions by GPs and compute their Maximum A Posteriori (MAP)/Maximum Likelihood Estimation (MLE) estimators given available data and constraints imposed by the structure of the graph. Taking into account the example depicted in Fig. 7, and substituting $f_{5,1}$, $f_{1,2}$, and $f_{3,6}$ with independent GPs, each with kernels $K,G$, and $\Gamma$ respectively, the objective of this MAP solution becomes minimizing $\|f_{5,1}\|_{K}^{2}+\|f_{1,2}\|_{G}^{2}+\|f_{3,6}\|_{\Gamma}^{2}$ (writing $\|f\|_{K}$ for the RKHS norm of $f$ induced by the kernel $K$) subject to the constraints imposed by the data and the functional dependencies encoded into the structure of the graph.
3.3. A system identification example.
In order to exemplify Computational Graphical Completion (CGC), consider the system identification problem depicted in Fig. 8, sourced from [29]. Our objective is to identify a nonlinear electric circuit, as illustrated in Fig. 8.(a), from scarce measurement data. The nonlinearity of the circuit emanates from the resistance, capacitance, and inductances, which are nonlinear functions of currents and voltages, as shown in Fig. 8.(b). Assuming these functions to be unknown, along with all currents and voltages as unknown time-dependent functions, we operate the circuit between times 0 and $10$. Measurements of a subset of variables, representing the system’s state, are taken at times $t_{s}=s/10$ for $s\in{0,\ldots,99}$. Given these measurements, the challenge arises in approximating all unknown functions that define currents and voltages as time functions, capacitance as a voltage function, and inductances and resistance as current functions. Fig. 8.(c) displays the available measurements, which are notably sparse, preventing us from reconstructing the underlying unknown functions independently. Thus, their interdependencies must be utilized for approximation. It is crucial to note that the system’s state variables are interconnected through functional relations, as per Kirchhoff’s laws for this nonlinear electric circuit, illustrated in Fig. 8.(d). These functional dependencies can be conceptualized as a computational graph, depicted in Fig. 8.(e), where nodes represent variables and directed edges represent functions. Known functions are colored in black, unknown functions in red, and round nodes aggregate variables, meaning edges map groups of variables, forming a hypergraph. The CGC solution involves substituting the graph’s unknown functions with Gaussian Processes (GPs), which may be independent or correlated, and then approximating the unknown functions with their Maximum A Posteriori (MAP) estimators, given the available data and the functional dependencies embedded in the graph’s structure. Fig. 8.(f) showcases the true and recovered functions, demonstrating a notably accurate approximation despite the data’s scarcity.
This simple example generalizes to an abstract framework detailed in [29]. This framework has a wide range of applications because most problems in CSE can also be formulated as completing computational graphs representing dependencies between functions and variables, and they can be solved in a similar manner by replacing unknown functions with GPs and by computing their MAP/EB estimator given the data. These problems include those illustrated in Fig. 1.(d-h).
4. Hardness and well-posed formulation of Type 3 problems.
@@ 3.2. Reminder on Computational Graph Completion for Type 2 problems
In this subsection, we describe why Type 3 problems are challenging and why they can even be intractable if not formalized and approached properly.
@@
Figure 9. Computational Hypergraph Discovery with three variables
4.2. Nonidentifiability and implicit dependencies.
Secondly, it is important to note that, even with an infinite amount of data, the exact structure of the hypergraph might not be identifiable. To illustrate this point, let’s consider a problem where we have $N$ samples from a computational graph with variables $x$ and $y$. The task is to determine the direction of functional dependency between $x$ and $y$. Does it go from $x$ to $y$ (represented as xxyyff ), or from $y$ to $x$ (represented as yyxxff )?
If we refer to Fig. 10.(a), we can make a decision because $y$ can only be expressed as a function of $x$. In contrast, if we examine Fig. 10.(b), the decision is also straightforward because $x$ can solely be written as a function of $y$. However, if the data mirrors the scenario in Fig. 10.(c), it becomes challenging to decide as we can write both $y$ as a function of $x$ and $x$ as a function of $y$. Further complicating matters is the possibility of implicit dependencies between variables. As illustrated in Fig. 10.(d), there might be instances where neither $y$ can be derived as a function of $x$, nor $x$ can be represented as a function of $y$.
@@
Causal inference methods broadly consist of two approaches: constraint and score-based methods. While constraint-based approaches are asymptotically consistent, they only learn the graph up to an equivalence class [44]. Instead, score-based methods resolve ambiguities in the graph’s edges by evaluating the likelihood of the observed data for each graphical model. For instance, they may assign a higher evidence to $y\to x$ over $x\to y$ if the conditional distribution $x|y$ exhibits less complexity than $y|x$. The complexity of searching over all possible graphs, however, grows super-exponentially with the number of variables. Thus, it is often necessary to use approximate, but more tractable, search-based methods [8, 34] or alternative criteria based on sensitivity analysis [10]. For example, the preference could lean towards $y\to x$ rather than $x\to y$ if $y$ demonstrates less sensitivity to errors or perturbations in $x$. In contrast, our proposed GP method avoids the growth in complexity by performing a guided pruning process that assesses the contribution of each node to the signal. We also emphasize that our method is not limited to learning acyclic graph structures as it can identify feedback loops between variables. Alternatively, methods for learning probabilistic undirected graphical models, also known as Markov networks, identify the graph structure by assuming the data is randomly drawn from some probability distribution [13]. In this case, edges in the graph (or lack thereof) encode conditional dependencies between the nodes. A common approach learns the graph structure by modeling the data as being drawn from a multivariate Gaussian distribution with a sparse inverse covariance matrix, whose zero entries indicate pairwise conditional independencies [14]. Recently, this approach has been extended using models for non-Gaussian distributions, e.g., in [1, 35], as well as kernel-based conditional independence tests [48]. In this work, we learn functional dependencies rather than causality or probabilistic dependence. We emphasize that we also do not assume the data is randomized or impose strong assumptions, such as additive noise models, in the data-generating process.
@@
We complete this paragraph by comparing the hypergraph discovery framework to structure learning for Bayesian networks and structural equation models (SEM). Let $x\in\mathbb{R}^{d}$ be a random variable with probability density function $p$ that follows the autoregressive factorization $p(x)=\prod_{i=1}^{d}p_{i}(x_{i}|x_{1},\dots,x_{i-1})$ given a prescribed variable ordering. Structure learning for Bayesian networks aims to find the ancestors of variable $x_{i}$, often referred to as the set of parents $Pa(i)\subseteq\{1,\dots,i-1\}$, in the sense that $p_{i}(x_{i}|x_{1},\dots,x_{i-1})=p_{i}(x_{i}|x_{Pa(i)})$. Thus, the variable dependence of the conditional density $p_{i}$ is identified by finding the parent set so that $x_{i}$ is conditionally independent of all remaining preceding variables given its parents, i.e., $x_{i}\perp x_{1:i-1\setminus Pa(i)}|x_{Pa(i)}$. Finding ancestors that satisfy this condition requires performing conditional independence tests, which are computationally expensive for general distributions [41]. Alternatively, SEMs assume that each variable $x_{i}$ is drawn as a function of its ancestors with additive noise, i.e, $x_{i}=f(x_{Pa(i)})+\epsilon_{i}$ for some function $f$ and noise $\epsilon$ [34]. For Gaussian noise $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, each marginal conditional distribution in a Bayesian network is given by $p_{i}(x_{i}|x_{1:i-1})\propto\exp(-\frac{1}{2\sigma^{2}}\|x_{i}-f(x_{1:i-1})\|^{2})$. Thus, finding the parents for such a model by maximum likelihood estimation corresponds to finding the parents that minimize the expected mean-squared error $\|x_{i}-f(x_{Pa(i)})\|^{2}.$ Our approach minimizes a related objective, without imposing the strong probabilistic assumptions that are required in SEMs and Bayesian Networks. We also observe that while the graph structure identified in Bayesian networks is influenced by the specific sequence in which variables are arranged (a concept exploited in numerical linear algebra [39, 37] where Schur complementation is equivalent to conditioning GPs and a carefully ordering leads to the accuracy of the Vecchia approximation $p_{i}(x_{i}|x_{1},\dots,x_{i-1})\approx p_{i}(x_{i}|x_{i-k},\dots,x_{i-1})$ [46]), the graph recovered by our approach remains unaffected by any predetermined ordering of those variables.
@@
Let $\mathcal{H}$ be a Reproducing Kernel Hilbert Space (RKHS) of functions mapping $\mathbb{R}^{d}$ to $\mathbb{R}$. Let $\mathcal{F}$ be a closed linear subspace of $\mathcal{H}$ and let $\mathcal{M}$ be a subset of $\mathbb{R}^{d}$ such that $x\in\mathcal{M}$ if and only if $f(x)=0$ for all $f\in\mathcal{F}$. Given the (possibly noisy and nonrandom) observation of $N$ elements, $X_{1},\ldots,X_{N}$, of $\mathcal{M}$ approximate $\mathcal{M}$.